# The Authentication Gap: Higher Education’s Widespread Noncompliance with NIST Digital Identity Guidelines

### v1

@@ 6 Discussion
Our study of organizational compliance with NIST standards could be framed and understood strictly in terms of incentives, resources, and priorities. From this perspective, the NIST standards constitute non-binding, technical guidance about authentication and digital identity that informs cost-benefit calculations, risk assessment/management, procurement, and other organizational decision making processes. Higher education organizations are complex, and various departments and personnel may be responsible for taking NIST standards into account. Resource constraints play a substantial role in cybersecurity, and as we noted above, variations in resources across institutions likely impacts (non)compliance with NIST standards. To increase the rate of compliance and consequently improve cybersecurity in higher education, society might need to re-engineer the relevant incentives, resources, and priorities. This might entail increasing public funding directed towards better cybersecurity practices (e.g., to support MFA), improving disclosure of policies and practices along the lines suggested previously (which might impact accountability to different stakeholders), and adjusting incentives and priorities through external pressure (e.g., insurance, markets, even legal reform).
However, this perspective does not provide a fully satisfactory explanation for the high rate of noncompliance with NIST standards that are quite easy and low cost to implement. Password expiration is the most obvious example. Refraining from requiring or recommending password replacement, except when a breach has occurred, is not difficult or expensive. Compliance is easy and should be widespread, yet our findings suggest it is not. We therefore believe the reason is not solely rooted in misaligned incentives, lack of resources, and alternative priorities. Instead, we think that this area is subject to a knowledge commons problem [19] concerning the diffusion of evolving expert knowledge, i.e., how such evolving knowledge translates (or not) into professional practice and system design.
The evolution of expert knowledge about authentication security over the course of decades has led to multiple 180-degree shifts in important NIST standards. The 2004 NIST SP 800-63 Electronic Authentication Guideline [8] advised users to protect and secure their accounts with complex passwords that comprise random characters, capital letters, and numbers, and to change their passwords regularly [29]. Federal agencies, corporations, and universities largely followed the advice by enforcing password complexity (e.g., mixtures of numbers, letters, and symbols), length, and/or regular password expiration. But expert knowledge on authentication systems changed dramatically as the field broadened beyond computational cybersecurity to include human factors, usability, and other knowledge sources [18]. Recognizing flaws and limitations of the 2004 publication, NIST published updates in 2011 (“800-63-1”), 2013 (“800-63-2”), and 2017 (“800-63-3”) [22]. In these successive updates, NIST established new guidelines for authentication to replace widespread practices that undermined security. One 180-degree revision, for example, was for organizations to no longer require periodic password expiration; instead, users should only change their passwords if there is evidence of compromise [21].
Frischmann and Johnson [18] suggest that our observed noncompliance with these updated recommendations may be attributable to knowledge commons dilemmas that inhibit diffusion of evolving expert security knowledge, initially to intermediate practitioners/professionals, such as software vendors and IT departments, and ultimately to laypeople. From this perspective, our study suggests that knowledge of changes in NIST standards may not have effectively diffused to many higher education institutions. If noncompliance is a consequence of delayed diffusion of knowledge, it may work itself out over time as more individuals gradually become aware of the updated wisdom. Alternatively, new forms of engagement may be necessary to facilitate the timely diffusion of expert knowledge reflected in NIST standards to those actually responsible for setting and implementing authentication policies and practices. This could involve continuing education and training courses at college and university IT departments, automated identification and notification of noncompliance by researchers or watchdog groups, or other to-be-developed methods of disseminating expert knowledge.
@@ 7 Conclusion
We examined online policies of 101 higher education institutions in the US and Canada to measure compliance with NIST SP 800-63-3 Digital Identity Guidelines. We focused on five standards that reflect changes from prior wisdom and impact all institutional account holders. We found broad deployment of multi-factor authentication, but widespread noncompliant use of password expiration, knowledge-based authentication, and password composition rules. These issues leave institutions vulnerable to cyberattacks, and less well-resourced institutions seem to have greater rates of noncompliance.
These results serve as a wake-up call that best practices that are well-understood by experts in the cybersecurity community have not sufficiently influenced the policies of higher education institutions. More investment and outreach in this area is needed, especially as improved authentication practices will harden institutions against the rising prevalence of ransomware and other cyberattacks.
@@ Acknowledgments
We thank Julia Tuck for her contributions. This research was supported by the Colgate University Faculty Research Council.

### v3

@@ 6 Discussion
The noncompliance we observe could be understood strictly in terms of incentives and resources. From this perspective, the NIST standards constitute non-binding, technical guidance about authentication and digital identity that informs cost-benefit calculations, risk assessment/management, procurement, and other organizational decision-making processes. Higher education organizations are complex, and various departments and personnel may be responsible for taking NIST standards into account. Resource constraints play a substantial role in cybersecurity [55], and variations in resources across institutions likely impact (non)compliance with NIST standards. Entrenched legacy systems [4], difficult to configure third-party software (e.g., single sign-on platforms [38]), staff shortages [22], and institutional prioritization of features other than cybersecurity could all contribute to resource shortages and incentives leading to noncompliance.
To increase the rate of compliance and consequently improve cybersecurity in higher education, society might need to re-engineer the availability of relevant resources and thereby change the corresponding incentives. This might entail increasing public funding directed towards better cybersecurity practices (e.g., to support MFA), improving disclosure of policies and practices (which might impact accountability to different stakeholders), or adjusting incentives and priorities through external pressure (e.g., insurance, markets, even legal reform).
However, this perspective does not provide a fully satisfactory explanation for persistent noncompliance with NIST standards that would otherwise reduce burden on users, decrease regular IT support costs, and improve authentication security. Password expiration is the most obvious example. Refraining from regular password replacement, except when a breach has occurred, reduces users’ time spent changing passwords and users’ mental effort remembering (or saving) new passwords. After initial system configuration updates, IT staff are relieved from recurring technical password expiration processes and from user support for lost or incorrect new passwords. Based on these aligned incentives, compliance should be widespread, yet our findings suggest it is not. This suggests that the noncompliance we observe is at least partially due to other factors, such as knowledge commons problems.
Explaining Noncompliance: Knowledge Diffusion.
Knowledge commons problems [32] concern the diffusion of evolving expert knowledge, i.e., how such evolving knowledge translates (or not) into professional practice and system design. Based primarily on interviews of cybersecurity experts, Frischmann and Johnson [31] have identified concerns that might impede expert knowledge diffusion, including that “professionals may be overconfident and inaccurately see themselves as security experts or as being sufficiently up-to-date on security.”
If the noncompliance we observe is indeed a consequence of a delayed diffusion of expert knowledge to many higher education institutions, it may work itself out over time as more practitioners gradually become aware of the updated standards. NIST itself has number of practices to facilitate the dissemination of its guidance, including public websites [59], direct engagement with universities and colleges [57], cybersecurity clinics [60], and mailing lists [58].
However, given the length of time since the NIST SP 800-63 Digital Identity Guidelines were published, our results suggest that new forms of engagement may be necessary to facilitate the timely diffusion of expert knowledge to those actually responsible for setting and implementing authentication policies and practices. There is an extensive usable security literature about the diffusion of software patches and other updates to users, system administrators, and network administrators [41, 50, 25, 49, 72] that could inform new methods of overcoming knowledge diffusion roadblocks in the cybersecurity standards space. Qualitative interviews of IT professionals at higher education institutions could also provide insight into overlooked or misunderstood reasons for persistent noncompliance.
We hope that the usable security community takes up this challenge, as a combination of human-computer interaction and sociotechnical systems research will be necessary to monitor and improve the rapid adoption of new best practices espoused in future standards updates.
@@ 7 Conclusion
We examined online policies of 135 higher education institutions in the US and Canada to measure compliance with NIST SP 800-63 Digital Identity Guidelines. We focused on four standards that reflect changes from prior versions and impact all institutional account holders. We found widespread, but not universal, deployment of multi-factor authentication. We also found widespread noncompliance with standards for password expiration, password composition rules, and knowledge-based authentication. These results indicate that while some expert cybersecurity recommendations have substantively influenced the policies of higher education institutions, more investment and outreach in this area is needed, as well as more research into incentive structures and the diffusion of expert guidance to practitioners.
@@ Acknowledgments
We thank Julia Tuck for her contributions. We also thank Max Alphonso, Pierce Anthony, Aly Bannister, William Beckhorn, Ryan Brackett, Matthew Burack, Onil Carrion, Jackson Carter, Sophia Cecchin, Danny Chu, Brian Douglas, Lily Ellis, Kate Gallagher, Julia Goosay, Jasper Gough, Alex Greene, Carly Grizzaffi, Nikki Izversky, Nephileen Kattel, Natalie McCall, Paige Mizutani, Aidan Murnane, Alex Tauber, Gray Theodore, and Blake Wang. This research was supported by the Colgate University Faculty Research Council. Yan Shvartzshnaider is supported in part by funding from the Social Sciences and Humanities Research Council of Canada.